If $20,000 is invested at 4% compound semiannually, how much is present at the end of 15 years? Round your answer to the nearest lowest dollar.
Thanks in advance :)

Answer by stanbon(57361)

If $20,000 is invested at 4% compound semiannually, how much is present at the end of 15 years? Round your answer to the nearest lowest dollar.
-----------------
A(t) = P(1+(r/n))^(nt)
-----
A(15) = 20,000*(1+(0.04/2))^(2*15)
--
A(15) = 20,000*(1.02)^30
A(15) = 20,000*1.811362
----
A(15) = $36227.23
======================
Cheers,
Stan H.
==============